Exponential equation without logs. Solve for x:
(4/5)^(-4x+5)= 64/125

Answer by stanbon(75887)

Exponential equation without logs. Solve for x:
(4/5)^(-4x+5)= 64/125
(4/5)^(-4x+5) = (4/5)^3
Since the bases are the same the exponents are equal:
-4x+5 = 3
-4x = -2
x = 1/2
============
Cheers,
Stan H.